OpenStudy (anonymous):

Check my work please?? Rate constant : 1.10x10^-4s^-1 @ 470C and 5.70x10^-4s^-1 @ 500C. What is the Ea?

OpenStudy (anonymous):

\[\ln \frac{ 5.70x10^-4 }{ 1.10x10^-4 }=-\frac{ Ea }{8.315x10^-3 }(\frac{ 1 }{ 773.15 }-\frac{ 1 }{ 743.15 })\]

OpenStudy (anonymous):

my answer = 2.69 kJ/ mol